OpenStudy (anonymous):

what percent of 49 is 7

hero (hero):

In general:\[\frac{\text{part}}{\text{whole}} = \frac{\%}{100}\] In this case \[\frac{7}{49} = \frac{x}{100}\] Solve for x to find the percent

OpenStudy (shamil98):

In addition to what hero displayed above. \[\large \frac{ 1 }{ 7 } = \frac{ x }{ 100 }\] You can reduce 7/49. From there you cross multiply the terms.